ICM 1998, Berlin, Aug. 18–27 Abstracts of Plenary and Invited Lectures

Section:

16. Applications

1991 MS Classification: 90C10 90C27 Cook, William, Rice University, USA The Traveling Salesman Problem (see Section 17)

ICM 1998 Berlin · Plenary and Invited Lectures